### v1

@@ 2.3 The Minimum Semantic Content (MSC) database: a reductionist approach to empirical aesthetics
In this study we build a novel database, the Minimum Semantic Content (MSC) database, which tackles the three issues that hamper a systematic analysis of the links between image features and aesthetic judgements, namely: (1) the entanglement between cognitive and perceptual valuation, (2) the systematic bias towards beautiful images, and (3) the lack of extreme examples. For this purpose, we first gather a large set of images that only contain objects from the natural environment (landscapes, trees, rocks, mountains, water, etc.) and therefore could be considered as lacking the elements that elicit the strongest cognitive and/or emotional responses, or, at least, to make semantic content uniform. Indeed, using “ugly” or synonyms as a keyword for retrieving internet images leads to outputs with strong emotional and cognitive content. For example, a quick Google search with the keywords “ugly images” produces a collection of grotesquely deformed human faces, suffering or diseased animals, shockingly violent images, feces, etc. This might be a consequence of humans evolving various emotional mechanisms such as fear, as a response to threats (e.g., predators, heights, starvation) or disgust (including sexual and moral disgust) as a response to risks and disease (Klebl2022, ). Moreover, studies on nature and urban imagery (Felisberti2022, ) suggest that the experience of ugliness is an independent experience not only entangled with the negation of beauty, but also able to coexist with it.
The new database we propose, characterized by a limited (arguably minimal) level of semantic content, enables us to explore the impact of image features on aesthetic judgement, free from the predominant influences of cognitive factors. In this regard, we adopt a reductionist methodology where the emotional and cognitive aspects of aesthetic judgments are held constant, facilitating the examination of the independent "perceptual" component (see Redies (Redies2015, )). Simultaneously, we mitigate the influence of multiple "themes" or "challenges" that could have contributed to the properties of the distribution in Figure 1.
@@
It is important to note that we explicitly avoided the inclusion of any type of artwork in our database. There is extensive literature showing the impact of contextual information on the aesthetic judgment of artworks (Darda2023, ), which we wanted to minimize. Similarly, we did not include databases of art or art-related images in our comparative analysis. We avoided art representations, as our goal was to create a controlled experimental setting that isolates perceptually-driven aesthetic decisions from those influenced by cognitive or semantic factors. By minimizing semantic content, the MSC dataset enables a focused examination of perceptual factors underlying aesthetic judgments without the confounding effects introduced by recognizable or culturally loaded content.
@@ 3.1 Dataset creation overview
  • Collect copyright free images. We collected a large amount of images exclusively from copyright-free or creative commons online databases (BurningWell2004, ; Flickr2003, ; GoodFreePhotos2012, ; Morguefile1996, ; PdPhoto2003, ; PhotosPublic2010, ) and own sources.
@@ 3.4 Obtaining aesthetic valuations from crowdsourcing results
After a visual inspection of the data, we decided to fit a truncated Gaussian (see Figure 4) to each histogram. This provided two statistical descriptors: the mean ($\mu$) and the SD of each valuation distribution. For the fittings, we assumed that the first bin corresponded to valuations between 0.5 to 1.5, the second bin to valuations between 1.5 and 2.5, and so on. Our fitting code was customized from Ryabov’s code (Ryabov2022, ) with the following constraints: (a) the value of $\mu$ was restricted between 0.5 and 5.5, which is the full range of the bars and (b) the SD of the Gaussian distribution was not allowed to have values below 0.5 (half of a histogram bar). Using these constraints, each image of the database was assigned a $\mu$ value in the range [0.5, 5.5] and a SD in the range [0.5, 1.9].
@@ 3.6 Aesthetic assessment metrics
Many image characteristics have been related to beauty, mainly from neuroscience and computational points of view (Brachmann2017, ; Datta2006, ). Since one of the objectives of this work is to showcase the importance of removing biases and semantics from the stimulus database, we selected a few good candidates features whose extraction is relatively easy to implement. Table 4, at the end of the section, provides a brief description of the metrics evaluated. For a justification, please refer to the discussion section below.
@@ 3.7 Basic, low-level features
  • Global contrast factor (GCF). This metric was obtained from the literature (Matkovic2005, ) and computes contrasts at various resolution levels in order to obtain a single measure of overall contrast. Weighting factors were obtained from psychophysical experiments in support of the authors’ claim that GCF better corresponds to human perception. A MATLAB version of the code was obtained from GitHub (Povoa2015, ).
@@
  • Overall perceived colorfulness. This metric was obtained from the literature (Hasler2003, ) and quantifies the (overall) perceived colorfulness of natural images from the distribution of the pixels in CIELab space using a linear combination of quantities (the SD along the a and b axis, the trigonometric length of the SD in ab space, the distance from the center of gravity in ab space to the neutral axis, etc.). The authors show that it correlates strongly with their psychophysical results. The MATLAB implementation was obtained from the GitHub repository (Povoa2015, ).
@@
  • Focus/Blur metric. This metric was also obtained from the literature (Thelen2009, ). It measures the relative degree of focus of an image using a diagonal Laplacian and was developed to obtain shape from focus for holographic reconstructions. Its effectiveness was compared to other up-to-date algorithms (Pertuz2013, ). The code was obtained from the MATLAB repository (Pertuz2022, ).
@@
  • Fourier amplitude slope ($\alpha$). This metric was chosen because of its links to physiological properties of the visual cortex (Field1987, ; Knill1990, ; Parraga2000, ), heightened sensitivity or preference toward natural spectra and visual discomfort (Penacchio2023, ; Penacchio2015, ). We calculated the Fourier amplitude slope in three steps: (a) convert all images to grayscales using MATLAB’s $rgb2gray$ function; (b) average Fourier amplitude over all orientations for each possible spatial frequency and, (c) plot the log of the resulting Fourier amplitude vs the logarithm of the spatial frequency and fit a line to the results. The slope of this line (a negative number) is the Fourier amplitude slope ($\alpha$). A higher negative value represents images with more Fourier energy in the low-spatial-frequencies (i.e., they contain large areas with relatively constant luminance), and a less negative value represents images with more Fourier energy in the high-spatial-frequencies (i.e., contain large areas with fine textures). The traditional way to measure $\alpha$ entails computing a value for every single spatial frequency component in Fourier space, averaging all orientations. When we plot these average Fourier amplitude values against spatial frequency in log-log space, most of the points are concentrated near the high end of the spectra, dominating the slope calculation. This problem becomes particularly bad in some images, e.g., those containing high energy noise. To counter the log-log space bias, we implemented another way of measuring $\alpha$: instead of averaging Fourier amplitude over all orientations for each possible spatial frequency, we averaged it for all orientations within spatial frequency “bands”. To achieve a logarithmically uniform distribution across frequencies, the widths of the bands were designed to increment in octaves. This improves line-fitting in log-log space, effectively reducing the impact of high spatial frequencies on the calculation of $\alpha$ (Brelstaff1995, ). We measured the slope of the amplitude spectrum using both methods (“traditional” and “unbiased”). It is important to notice that a steeper $\alpha$ correlates with, but is not the same as blurring, since it involves changes at both ends of the Fourier spectrum.
@@
  • Image complexity. This consists of two metrics developed by Groen et al (Groen2013, ) to describe the complexity of an image in terms of its contrast distribution and scene fragmentation using summary statistics obtained from the spatial pooling of responses of neurons in the visual pathway of the brain (Scholte2009, ). The first metric, called Contrast Energy (CE), describes the scale of contrast in an image. The second metric, called Spatial Coherence (SC), describes the degree of correlation between contrast elements in the image. Both metrics convolve the image with multiscale filters (smaller sizes for CE than for SC) and for each image location, a single filter response is selected and pooled across a selection of the visual field (for CE responses were averaged and for SC a coefficient of variation was obtained). These were later averaged across the three color-opponent representations resulting in one CE and SC value per image. In their article, Groen et al. show that CE describes the scale of the contrast distribution (varies with the distribution of local contrasts strengths) and SC describes the shape of the contrast distribution (varies with the amount of scene fragmentation/clutter). They also show that these metrics (SC in particular) are related to single-trial event-related potentials (sERPs) obtained during a naturalness rating task. The code to compute CE and SC was kindly provided by Iris Groen. We calculated both metrics using the default parameters for their code.
@@ 3.8 Higher level features (machine-learning-based)
  • Symmetry. We tried the symmetry metric developed by Brachmann and Redies (Brachmann2016, ) who used a modification of CaffeNet (Krizhevsky2012, ), a Convolutional Neural Network (CNN) that was trained on the ImageNet database. They validated their algorithm on music album covers, which were rated according to their symmetry by human observers. The algorithm provides several options, and we chose the Left-Right and Top-Down variants. The authors claim that their model of symmetry closely resembles human perception of symmetry in CD music album covers. A Python version of the code was obtained from OSF (Redies2019, ).
  • Mean 3D depth. This monocular metric (Ranftl2021, ) tries to quantify the mean depth of the image using dense prediction transformers, an architecture that leverages vision transformers instead of convolutional networks. Images in the database were reduced to have a maximum side size of 150 pixels to save computer processing power. The PyTorch-converted models were obtained from GitHub (Ranftl2012, ).
@@
Global Contrast Factor (Matkovic2005, ) Single measure of overall contrast, psychophysically tested.
@@
Perceived Colorfulness (Hasler2003, ) Linear combination of statistical quantities in CIELab, psychophysically tested.
@@
Focus/Blur metric (Pertuz2013, ) Relative degree of focus of an image using a diagonal Laplacian.
@@
Image Complexity (Groen2013, ) Metrics describing the complexity of an image in terms of its contrast energy distribution and spatial coherence (scene fragmentation).
@@
L-R and U-D Symmetry metric (Brachmann2016, ) Left-Right and Up-Down variants of a symmetry metric obtained by a CNN trained on the ImageNet database. Validated by human observers on CD album covers.
Mean 3D-depth (Ranftl2021, ) Quantifies the mean depth of the image using dense prediction transformers.
@@ 3.9 Results
We applied the above metrics to the MSC database and obtained correlations with the mean ratings $\mu$. Given that auto-uglified images were the product of applying the uglification in a random manner and were mostly considered ugly in the crowdsourcing (see Figure 5), we combined them with the uglified for our analysis. Table 5 summarizes the correlation results including an extra column for those of AVA (Murray2012, ), which is arguably one of the best documented and popular databases available today. To compare MSC to AVA we split our correlation plots into two categories: all images and unmodified images only. The first allows us to visualize the contribution of modified images to the results and the second underlines how, by simply removing semantics, we can unveil correlations that are hidden in other databases. In summary, the difference between the first and third columns in Table 5 highlights the impact of removing the bias towards beautiful imagery, and the difference between the third and fifth columns highlights the confounding role of semantics in aesthetics research.
Detailed results for each metric applied to MSC are presented below. We separated the MSC into three categories of images: (a) uglified & auto-uglified, (b) unmodified and (c) beautified. Because of the large number of datapoints, we highlighted prominent clusters using colored scatterplots from the MATLAB repository (Lukas2023, ). Correlations (r) refer to Pearson’s r and their corresponding p-values. Also because of the large number of datapoints, many p-values are extremely small (e.g., p< 10-10) and are quoted as “p<0.001”. The exact p-values are in Table 5.
@@
Global Contrast Factor (Matkovic2005, ) 0.220 $10^{-104}$0.004 0.7 0.004 0.03
@@
Focus& Blur metric (Pertuz2013, ) -0.037 0.0001 0.025 0.05 0.064 $10^{-228}$
@@
Image Complexity (Groen2013, ) 0.171 0 -0.002 0.877 0.043 $10^{-105}$
L-R Symmetry metric (Brachmann2016, ) -0.344 $10^{-288}$0.270 $10^{-95}$0.026 $10^{-42}$
U-D Symmetry metric (Brachmann2016, ) -0.361 $10^{-317}$0.337 $10^{-150}$0.008 0
Mean 3D-depth (Ranftl2021, ) -0.188 $10^{-84}$0.271 $10^{-96}$0.085 0
@@ 3.13 Overall perceived colorfulness, and saturation from the HSV color model
Two other metrics related to color, overall colorfulness (Hasler2003, ) and saturation from the HSV color model, produced very similar results to those of Figure 8. In the case of overall colorfulness, the correlation was negligible ($r=-0.029$, $p<0.001$) for all images, and positive and significant ($r=0.356$, $p<0.001$) for the subset of unmodified images. In the case of saturation, the correlation was negligible ($r=-0.088$, $p<0.001$) for all images and somewhat larger ($r=0.258$, $p<0.001$) for unmodified images. Again, modified, low-$\mu$ images populated the full range of colorfulness on the left side of picture A, in effect making the correlation value negligible. These results are shown in the appendix (Figure A5 and Figure A6).
@@ 3.16 Fourier amplitude slope ($\alpha$)
Figure 11: Results for the Image Complexity metric (Groen2013, ) considering only Contrast Energy. (A) Scatterplot for all images. (B) Same for unmodified images. (C) Metric values for the three groups of images considered. Boxplots as in previous figures.

### v2

@@ 2.3 The Minimum Semantic Content (MSC) database: a reductionist approach to empirical aesthetics
In this study we build a novel database, the Minimum Semantic Content (MSC) database, which tackles the three issues that hamper a systematic analysis of the links between image features and aesthetic judgements, namely: (1) the entanglement between cognitive and perceptual valuation, (2) the systematic bias towards beautiful images, and (3) the lack of extreme examples. For this purpose, we first gather a large set of images that only contain objects from the natural environment (landscapes, trees, rocks, mountains, water, etc.) and therefore could be considered as lacking the elements that elicit the strongest cognitive and/or emotional responses, or, at least, to make semantic content uniform. Indeed, using “ugly” or synonyms as a keyword for retrieving internet images leads to outputs with strong emotional and cognitive content. For example, a quick Google search with the keywords “ugly images” produces a collection of grotesquely deformed human faces, suffering or diseased animals, shockingly violent images, feces, etc. This might be a consequence of humans evolving various emotional mechanisms such as fear, as a response to threats (e.g., predators, heights, starvation) or disgust (including sexual and moral disgust) as a response to risks and disease [71]. Moreover, studies on nature and urban imagery [72] suggest that the experience of ugliness is an independent experience not only entangled with the negation of beauty, but also able to coexist with it.
The new database we propose, characterized by a limited (arguably minimal) level of semantic content, enables us to explore the impact of image features on aesthetic judgement, free from the predominant influences of cognitive factors. In this regard, we adopt a reductionist methodology where the emotional and cognitive aspects of aesthetic judgments are held constant, facilitating the examination of the independent "perceptual" component (see Redies [73]). Simultaneously, we mitigate the influence of multiple "themes" or "challenges" that could have contributed to the properties of the distribution in Figure 1.
@@
It is important to note that we explicitly avoided the inclusion of any type of artwork in our database. There is extensive literature showing the impact of contextual information on the aesthetic judgment of artworks [74], which we wanted to minimize. Similarly, we did not include databases of art or art-related images in our comparative analysis. We avoided art representations, as our goal was to create a controlled experimental setting that isolates perceptually-driven aesthetic decisions from those influenced by cognitive or semantic factors. By minimizing semantic content, the MSC dataset enables a focused examination of perceptual factors underlying aesthetic judgments without the confounding effects introduced by recognizable or culturally loaded content.
@@ 3.1 Dataset creation overview
  • Collect copyright free images. We collected a large amount of images exclusively from copyright-free or creative commons online databases [75, 56, 76, 77, 78, 79] and own sources.
@@ 3.4 Obtaining aesthetic valuations from crowdsourcing results
After a visual inspection of the data, we decided to fit a truncated Gaussian (see Figure 4) to each histogram. This provided two statistical descriptors: the mean ($\mu$) and the SD of each valuation distribution. For the fittings, we assumed that the first bin corresponded to valuations between 0.5 to 1.5, the second bin to valuations between 1.5 and 2.5, and so on. Our fitting code was customized from Ryabov’s code [80] with the following constraints: (a) the value of $\mu$ was restricted between 0.5 and 5.5, which is the full range of the bars and (b) the SD of the Gaussian distribution was not allowed to have values below 0.5 (half of a histogram bar). Using these constraints, each image of the database was assigned a $\mu$ value in the range [0.5, 5.5] and a SD in the range [0.5, 1.9].
@@ 3.6 Aesthetic assessment metrics
Many image characteristics have been related to beauty, mainly from neuroscience and computational points of view [81, 50]. Since one of the objectives of this work is to showcase the importance of removing biases and semantics from the stimulus database, we selected a few good candidates features whose extraction is relatively easy to implement. Table 4, at the end of the section, provides a brief description of the metrics evaluated. For a justification, please refer to the discussion section below.
@@ 3.7 Basic, low-level features
  • Global contrast factor (GCF). This metric was obtained from the literature [82] and computes contrasts at various resolution levels in order to obtain a single measure of overall contrast. Weighting factors were obtained from psychophysical experiments in support of the authors’ claim that GCF better corresponds to human perception. A MATLAB version of the code was obtained from GitHub [83].
@@
  • Overall perceived colorfulness. This metric was obtained from the literature [84] and quantifies the (overall) perceived colorfulness of natural images from the distribution of the pixels in CIELab space using a linear combination of quantities (the SD along the a and b axis, the trigonometric length of the SD in ab space, the distance from the center of gravity in ab space to the neutral axis, etc.). The authors show that it correlates strongly with their psychophysical results. The MATLAB implementation was obtained from the GitHub repository [85].
@@
  • Focus/Blur metric. This metric was also obtained from the literature [86]. It measures the relative degree of focus of an image using a diagonal Laplacian and was developed to obtain shape from focus for holographic reconstructions. Its effectiveness was compared to other up-to-date algorithms [87]. The code was obtained from the MATLAB repository [88].
@@
  • Fourier amplitude slope ($\alpha$). This metric was chosen because of its links to physiological properties of the visual cortex [89, 90, 91], heightened sensitivity or preference toward natural spectra and visual discomfort [92, 93]. We calculated the Fourier amplitude slope in three steps: (a) convert all images to grayscales using MATLAB’s $rgb2gray$ function; (b) average Fourier amplitude over all orientations for each possible spatial frequency and, (c) plot the log of the resulting Fourier amplitude vs the logarithm of the spatial frequency and fit a line to the results. The slope of this line (a negative number) is the Fourier amplitude slope ($\alpha$). A higher negative value represents images with more Fourier energy in the low-spatial-frequencies (i.e., they contain large areas with relatively constant luminance), and a less negative value represents images with more Fourier energy in the high-spatial-frequencies (i.e., contain large areas with fine textures). The traditional way to measure $\alpha$ entails computing a value for every single spatial frequency component in Fourier space, averaging all orientations. When we plot these average Fourier amplitude values against spatial frequency in log-log space, most of the points are concentrated near the high end of the spectra, dominating the slope calculation. This problem becomes particularly bad in some images, e.g., those containing high energy noise. To counter the log-log space bias, we implemented another way of measuring $\alpha$: instead of averaging Fourier amplitude over all orientations for each possible spatial frequency, we averaged it for all orientations within spatial frequency “bands”. To achieve a logarithmically uniform distribution across frequencies, the widths of the bands were designed to increment in octaves. This improves line-fitting in log-log space, effectively reducing the impact of high spatial frequencies on the calculation of $\alpha$ [94]. We measured the slope of the amplitude spectrum using both methods (“traditional” and “unbiased”). It is important to notice that a steeper $\alpha$ correlates with, but is not the same as blurring, since it involves changes at both ends of the Fourier spectrum.
@@
  • Image complexity. This consists of two metrics developed by Groen et al [95] to describe the complexity of an image in terms of its contrast distribution and scene fragmentation using summary statistics obtained from the spatial pooling of responses of neurons in the visual pathway of the brain [96]. The first metric, called Contrast Energy (CE), describes the scale of contrast in an image. The second metric, called Spatial Coherence (SC), describes the degree of correlation between contrast elements in the image. Both metrics convolve the image with multiscale filters (smaller sizes for CE than for SC) and for each image location, a single filter response is selected and pooled across a selection of the visual field (for CE responses were averaged and for SC a coefficient of variation was obtained). These were later averaged across the three color-opponent representations resulting in one CE and SC value per image. In their article, Groen et al. show that CE describes the scale of the contrast distribution (varies with the distribution of local contrasts strengths) and SC describes the shape of the contrast distribution (varies with the amount of scene fragmentation/clutter). They also show that these metrics (SC in particular) are related to single-trial event-related potentials (sERPs) obtained during a naturalness rating task. The code to compute CE and SC was kindly provided by Iris Groen. We calculated both metrics using the default parameters for their code.
@@ 3.8 Higher level features (machine-learning-based)
  • Symmetry. We tried the symmetry metric developed by Brachmann and Redies [97] who used a modification of CaffeNet [98], a Convolutional Neural Network (CNN) that was trained on the ImageNet database. They validated their algorithm on music album covers, which were rated according to their symmetry by human observers. The algorithm provides several options, and we chose the Left-Right and Top-Down variants. The authors claim that their model of symmetry closely resembles human perception of symmetry in CD music album covers. A Python version of the code was obtained from OSF [99].
  • Mean 3D depth. This monocular metric [100] tries to quantify the mean depth of the image using dense prediction transformers, an architecture that leverages vision transformers instead of convolutional networks. Images in the database were reduced to have a maximum side size of 150 pixels to save computer processing power. The PyTorch-converted models were obtained from GitHub [101].
@@
Global Contrast Factor [82] Single measure of overall contrast, psychophysically tested.
@@
Perceived Colorfulness [84] Linear combination of statistical quantities in CIELab, psychophysically tested.
@@
Focus/Blur metric [87] Relative degree of focus of an image using a diagonal Laplacian.
@@
Image Complexity [95] Metrics describing the complexity of an image in terms of its contrast energy distribution and spatial coherence (scene fragmentation).
@@
L-R and U-D Symmetry metric [97] Left-Right and Up-Down variants of a symmetry metric obtained by a CNN trained on the ImageNet database. Validated by human observers on CD album covers.
Mean 3D-depth [100] Quantifies the mean depth of the image using dense prediction transformers.
@@ 3.9 Results
We applied the above metrics to the MSC database and obtained correlations with the mean ratings $\mu$. Given that auto-uglified images were the product of applying the uglification in a random manner and were mostly considered ugly in the crowdsourcing (see Figure 5), we combined them with the uglified for our analysis. Table 5 summarizes the correlation results including an extra column for those of AVA [54], which is arguably one of the best documented and popular databases available today. To compare MSC to AVA we split our correlation plots into two categories: all images and unmodified images only. The first allows us to visualize the contribution of modified images to the results and the second underlines how, by simply removing semantics, we can unveil correlations that are hidden in other databases. In summary, the difference between the first and third columns in Table 5 highlights the impact of removing the bias towards beautiful imagery, and the difference between the third and fifth columns highlights the confounding role of semantics in aesthetics research.
Detailed results for each metric applied to MSC are presented below. We separated the MSC into three categories of images: (a) uglified & auto-uglified, (b) unmodified and (c) beautified. Because of the large number of datapoints, we highlighted prominent clusters using colored scatterplots from the MATLAB repository [102]. Correlations (r) refer to Pearson’s r and their corresponding p-values. Also because of the large number of datapoints, many p-values are extremely small (e.g., p< 10-10) and are quoted as “p<0.001”. The exact p-values are in Table 5.
@@
Global Contrast Factor [82] 0.220 $10^{-104}$0.004 0.7 0.004 0.03
@@
Focus& Blur metric [87] -0.037 0.0001 0.025 0.05 0.064 $10^{-228}$
@@
Image Complexity [95] 0.171 0 -0.002 0.877 0.043 $10^{-105}$
L-R Symmetry metric [97] -0.344 $10^{-288}$0.270 $10^{-95}$0.026 $10^{-42}$
U-D Symmetry metric [97] -0.361 $10^{-317}$0.337 $10^{-150}$0.008 0
Mean 3D-depth [100] -0.188 $10^{-84}$0.271 $10^{-96}$0.085 0
@@ 3.13 Overall perceived colorfulness, and saturation from the HSV color model
Two other metrics related to color, overall colorfulness [84] and saturation from the HSV color model, produced very similar results to those of Figure 8. In the case of overall colorfulness, the correlation was negligible ($r=-0.029$, $p<0.001$) for all images, and positive and significant ($r=0.356$, $p<0.001$) for the subset of unmodified images. In the case of saturation, the correlation was negligible ($r=-0.088$, $p<0.001$) for all images and somewhat larger ($r=0.258$, $p<0.001$) for unmodified images. Again, modified, low-$\mu$ images populated the full range of colorfulness on the left side of picture A, in effect making the correlation value negligible. These results are shown in the appendix (Figure A5 and Figure A6).
@@ 3.16 Fourier amplitude slope ($\alpha$)
Figure 11: Results for the Image Complexity metric [95] considering only Contrast Energy. (A) Scatterplot for all images. (B) Same for unmodified images. (C) Metric values for the three groups of images considered. Boxplots as in previous figures.